how many gallons of 20% alcohol soultion and 50% alcohol soultion must be mixed together to get 9 gallons of 30% alcohol solution?

Answer by mananth(16946)

percent ---------------- quantity
Alcohol I 20 ---------------- x gallons
Alcohol II 50 ------ 9 - x gallons
Mixture 30.00% ---------------- 9

20x+50( 9-x)=30*9

20x+450 -50x= 270
20x-50x =270-450
-30x=-180
/-30
x = 6 gallons 20.00% Alcohol I
3 gallons 50.00% Alcohol II

m.ananth@hotmail.ca